Assume that the average firm in your company's industry is expected to grow at a constant rate of 4% and that its dividend yield is 8%. Your company is about as risky as the average firm in the industry and just paid a dividend (D0) of $2. You expect that the growth rate of dividends will be 50% during the first year (g0,1 = 50%) and 30% during the second year (g1,2 = 30%). After Year 2, dividend growth will be constant at 4%. What is the estimated value per share of your firm’s stock? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

Please find below solution..

Required rate = Growth rate + Dividend yield rate
Required rate = =8%+4% 12.00%
Value of stock = present value of future cash flow
Dividend year 1 = =2*(1+50%) 3.00
Dividend year 2 = =3*(1+30%) 3.90
Dividend year 3 = =3.90*(1+4%) 4.056
Computation of horizon value = Dividend in year 3/ (required rate - growth rate)
=4.056/(12%-4%)
50.70
Computation of share price
i ii iii=i+ii iv v=iv*iii
Year Dividend Terminal value Total cash flow PVIF 12% present value
1 3.00 3.00     0.8929          2.68
2 3.90 50.70 54.60     0.7972       43.53
      46.21
Price today =                 46.21